OpenStudy (anonymous):

Urgent help needed: Expand the following up to the term in x^3:

OpenStudy (anonymous):

\[\frac{ 1+x }{ \sqrt{1-x} }\]

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

expand in a power series i take it

OpenStudy (anonymous):

about zero i hope first term is \(f(0)=1\)

OpenStudy (anonymous):

first derivative is \[f'(x)=\frac{3-x}{2\sqrt{1-x}^3}\] so second term is \[f'(0)x=\frac{3}{2}x\]

OpenStudy (anonymous):

yikes, now we need the derivative of that beast

OpenStudy (anonymous):

do i have to do that too?

OpenStudy (anonymous):

Yes, please?

OpenStudy (anonymous):

grrr

OpenStudy (anonymous):

lets cheat

OpenStudy (anonymous):

looks like it is \[f''(x)=\frac{7-x}{4\sqrt{1-x}^5}\]

OpenStudy (anonymous):

making the next term \[\frac{f''(0)}{2}x^2=\frac{7}{8}x^2\]

OpenStudy (anonymous):

one more?

OpenStudy (anonymous):

So I find the derivative of that too?

OpenStudy (anonymous):

\[f'^{(3)}(x)=\frac{33-3x}{8\sqrt{1-x}^7}\]

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

if so then \(\frac{f^{(3)}(0)}{3!}x^3=\frac{33}{16}x^3\)

OpenStudy (anonymous):

hmm let me check if this is right

OpenStudy (anonymous):

no messed up \(3!=6\) doh

OpenStudy (anonymous):

third term is \(\frac{11}{16}x^3\)

OpenStudy (anonymous):

k lets check the whole dam thing and see if it is right

OpenStudy (anonymous):

Okay, thank you so much! :)

OpenStudy (anonymous):

yw